OpenStudy (anonymous):

(32 − 32i)^-6 how would i write this in polar form?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

I guess you first have to write \(32-32i\) in polar form right?

OpenStudy (misty1212):

any idea how to do that? it is not too hard, you need two numbers \(r\) and \(\theta\)

OpenStudy (anonymous):

i was wondering if my r would be \[\sqrt{2048}\]

OpenStudy (anonymous):

and how i would turn that in \[\theta \]

OpenStudy (anonymous):

wouldnt my theta be \[32/\sqrt{2048}+-32/\sqrt{2048}i\]

OpenStudy (anonymous):

and turn that into cos and sin

OpenStudy (anonymous):

@misty1212

zepdrix (zepdrix):

Hey Art :) Here is a neat little trick, it's probably not the way you're used to solving these though.\[\Large\rm \left(32-32\mathcal i\right)^{-6}\]Whenever I see that the numerical values are the same for the real and imaginary part, I immediately know that this is some type of pi/4 angle, because sine and cosine give us the same value (maybe different sign) at pi/4, ya? So what I want to do is, try to turn these 32's into sqrt2/2's. Because then I can work backwards and turn my sqrt2/2's into some kind of cosine and sine of one of the pi/4 angles. So I'll factor the 32 out of each term, and also another 2, so we end up with a 2 in the bottom. So a 64 in total.\[\Large\rm =64^{-6}\left(\frac{1}{2}-\frac{1}{2}\mathcal i\right)^{-6}\]I have to remember to apply my -6 whenever I pull something out. I want to `gain` a sqrt2 in the top, so I'll divide out a \(\Large\rm \frac{1}{\sqrt2}\).\[\Large\rm =\frac{64^{-6}}{\sqrt2^{-6}}\left(\frac{\sqrt2}{2}-\frac{\sqrt2}{2}\mathcal i\right)^{-6}\]Let's get rid of the negative in that exponent out front,\[\Large\rm =\left(\frac{\sqrt2}{64}\right)^6\left(\color{orangered}{\frac{\sqrt2}{2}-\frac{\sqrt2}{2}\mathcal i}\right)^{-6}\]So umm, for this orange part... Cosine corresponds to the real part, sine the imaginary part, ya? So we want to be in the quadrant where: `cosine is positive` `sine is negative` Puts in quadrant 4. So looks like we want 7pi/4.

zepdrix (zepdrix):

\[\Large\rm =\left(\frac{\sqrt2}{64}\right)^6\left(\color{orangered}{\cos\left(\frac{7\pi}{4}\right)+\mathcal i\sin\left(\frac{7\pi}{4}\right)}\right)^{-6}\]

zepdrix (zepdrix):

Then of course from here we can apply De'Moivre's Theorem! :)\[\Large\rm =\left(\frac{\sqrt2}{64}\right)^6\left(\color{orangered}{\cos\left(\frac{-6\cdot7\pi}{4}\right)+\mathcal i\sin\left(\frac{-6\cdot7\pi}{4}\right)}\right)\]

zepdrix (zepdrix):

From there you just need to simplify your radius a bit, and simplify your angle, unwind it.

zepdrix (zepdrix):

Not the most typical way of doing this, I apologize if it's too confusing :)

OpenStudy (anonymous):

I understand somewhat but I see how you got everything and from there ill put the pieces together but that was a pretty sweet explanation. Thanks!!!!!